# Benchmark the Linguistic Competence of Language Models

### v1

@@ 6 Efficiency
Besides excluding licensed data (18 probing datasets), we analyze the effect of discarding training instances. As a result, we reduce the computation for encoding and the actual probing simultaneously. We follow Perlitz et al. (2023) and calculate the rank resolution, $95\%$ CI of model rank difference. This measure indicates the maximum expected rank deviation from evaluating an LM on FlashHolmes compared to Holmes. For example, a rank resolution of one means that an LM evaluated on FlashHolmes and Holmes has the same rank or switch place with its neighbors with a probability of 95%. Figure 12 shows the resulting rank resolution when training only on a fraction of the instances, from $1/2$ to $1/512$. Solely focusing on efficiency ($1/512$) still provides a decent rank resolution of ~2.7. In contrast, considering $1/2$ of the training data results in the best reliability of ~1.0. To balance benchmark reliability and efficiency, we compose FlashHolmes using $1/32$ of the training instances. Precisely, it reduces the computation expenses of evaluating LMs to ~3% of what Holmes would have required while preserving a high rank-correlation of ~1.3.
@@ 7 Related Work
Benchmarks approximate LMs abilities like general language understanding (Wang et al., 2019b, a), out-of-distribution generalization (Yang et al., 2023; Waldis et al., 2024b), adversarial scenarios (Nie et al., 2020; Wang et al., 2021), or retrieval like BEIR (Thakur et al., 2021) or MTEB (Muennighoff et al., 2023). With the advent of larger LMs, the methodological focus shifted to prompting-based evaluations which evaluate the LMs’ response to provided instructions (Brown et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022) covering application-oriented tasks (Liang et al., 2023), or mathematical reasoning (e.g., GSM8K Cobbe et al. (2021)).
@@
The analysis of LMs’ linguistic competence ranges from analyzing static word vectors (Köhn, 2015), sentence embeddings (Conneau et al., 2018; Adi et al., 2017), the internals of translation models (Shi et al., 2016; Bau et al., 2019), or contextualized LMs (Tenney et al., 2019b, a; Hewitt and Manning, 2019). Other work addressed methodological aspects, such as using control tasks (Hewitt and Liang, 2019), assessing LMs from an information theory perspective (Voita and Titov, 2020; Pimentel et al., 2020), or evaluating causal effects in LMs (Elazar et al., 2021). Finally, another line of work focuses on whether LMs follow human understanding of linguistic competence when solving downstream tasks (Belinkov, 2022; Aw et al., 2023; Mahowald et al., 2024). However, Mosbach et al. (2020b) and Waldis et al. (2024a) found fine-tuning for downstream tasks actually hurting the understanding of linguistic phenomena.
While prior studies assessing the linguistic competence of LMs tend to focus on a limited set of linguistic phenomena or models, Holmes provides extensive coverage of both phenomena and evaluated LMs. Unlike recent evaluations based on prompting methods (Blevins et al., 2023; Liang et al., 2023; Amouyal et al., 2024), Holmes assesses the internal representations of LMs directly. This approach allows for detailed analysis of specific model characteristics, such as architecture, and helps separate the linguistic competence from other cognitive abilities. Thereby, we respond to recent calls for a thorough and explicit evaluation of linguistic phenomena (Hu and Levy, 2023; Lu et al., 2023; Mahowald et al., 2024).
@@ 8 Conclusion
Holmes marks the most up-to-date and extensive consolidation of existing resources addressing the need to assess the linguistic competence of LMs in isolation. Our experiments demonstrate that LMs’ linguistic competence is pronounced regarding formal phenomena but lacks functional ones when information about broader textual contexts, such as rhetorical structure, is required. Further, size, architecture, and instruction tuning crucially account for differences among LMs. As LM and resources in the landscape of linguistics continue to grow, we will actively extend Holmes with further probing datasets, evaluate upcoming LMs, and plan to incorporate multilingualism.
@@
Phenomena and LM Coverage
We agree with Liang et al. (2023) and see one fundamental aspect in composing a benchmark in acknowledging its incompleteness. Both linguistic phenomena and LMs are a subset of the variety of available resources. We consolidated them carefully to provide a comprehensive scope of the linguistic competence and various LMs. However, as benchmarks evolve as tools to assess LMs, we will further expand Holmes both with the existing and upcoming LMs and data resources.
@@
As Holmes relies on existing resources, it inherits the bias embodied in this data. Examples of such bias are gender equality or gender fairness, like the use of neo pronouns such as em in Lauscher et al. (2023).
@@ A.1 Additional Details on the Evolution of Probing Literature
We analyze publication trends by year and venue as shown in Table 2. Less work was published between 2015-2018 (earlier) focusing on LSTM-based (Linzen et al., 2016; Conneau et al., 2018) and static LMs (Köhn, 2015; Linzen et al., 2016; Belinkov et al., 2017; Conneau et al., 2018). With the release of BERT (Devlin et al., 2019) in 2019, we note increasing attention to analyzing linguistic abilities within LMs, with a peak of 90 papers in 2022.444Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-analysis. Considering the venue, more than half of the relevant work (149 papers) was published at major conferences (ACL and EMNLP), and 68 papers were published at AACL, EACL, NAACL, and COLING. [^5] Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-study. In addition, we observe a constant contribution of TACL, various workshops, such as Analyzing and Interpreting Neural Networks for NLP or Representation Learning for NLP.
@@ A.2 Experimental Details
A.3 Linguistic Task Categorization
We show in Table 3, Table 4, Table 7, Table 5, and Table 6 which resources Holmes use to cover morphology, syntax, semantics, reasoning, and discourse phenomena. This includes 33 works providing the data, the specific linguistic phenomena, or both. For example, for readability we use the data of Weischedel et al. (2013) and calculated the flesch score (Flesch, 1948).
@@ A.3 Linguistic Task Categorization
The second group of 75 tasks verifies the following syntax phenomena: Part-of-speech tagging and constituent labeling (Weischedel et al., 2013); dependency labeling (Silveira et al., 2014); bigram-shift, tree-depth, top-constituent-task, and sentence-length (Conneau et al., 2018); subject- & object-number, and deoncausative-inchoative alternation based on Klafka and Ettinger (2020); binding, control/raising, negative polarity item licensing, island-effects, argument-structure, ellipsis, and filler-gap (Warstadt et al., 2020; Huebner et al., 2021).
@@
Third, consider 67 tasks covering semantics phenomena: Named-entity labeling and semantic-role labeling (Weischedel et al., 2013); subject- and object-number, tense, semantic odd man out, word content, and coordination inversion (Conneau et al., 2018); semantic relation classification (Hendrickx et al., 2010); semantic proto-roles (Rudinger et al., 2018a); factuality (Rudinger et al., 2018b); genericity (Govindarajan et al., 2019); event structure (Gantt et al., 2022); time (Vashishtha et al., 2019); word sense (White et al., 2016); sentiment analysis (Socher et al., 2013); object- and subject-animacy, object- and subject-gender, verb-tense, and verb-dynamic Klafka and Ettinger (2020); metaphor (Mohler et al., 2016; Birke and Sarkar, 2006; Steen et al., 2010); complex word identification (Paetzold and Specia, 2016); and passive (Krasnowska-Kieraś and Wróblewska, 2019). In addition, we derive synonym-/antonym-detection task using WordNet (Miller, 1995) and the texts from OntoNotes v5 Weischedel et al. (2013).
@@
Forth, 19 tasks cover reasoning phenomena: Paraphrasticity with negation and antonyms (Vahtola et al., 2022); negation detection (Szarvas et al., 2008; Konstantinova et al., 2012; Morante and Blanco, 2012); negation-span classification (Szarvas et al., 2008; Konstantinova et al., 2012); negation-correspondence (Szarvas et al., 2008; Konstantinova et al., 2012); speculation detection, speculation-span classification, and speculation-correspondence (Szarvas et al., 2008); and always-never, age comparison, objects comparison, antonym negation, property conjunction, taxonomy connection, encyclopedic composition, and multi-hop composition (Talmor et al., 2020).
@@
Finally, Holmes embodies 28 task addressing discourse phenomena: Co-reference resolution Weischedel et al. (2013); bridging (Hou, 2018, 2020; Pandit and Hou, 2021); discourse connective (Nie et al., 2019); sentence order and next-sentence prediction (Narayan et al., 2018); discourse correspondence, discourse order, discourse relation, discourse distance, discourse explicit classes, discourse implicit classes (Webber et al., 2019; Kurfalı and Östling, 2021); and rst-count/-depth/-distance/-relation/-relation-group/-successively/-type (Carlson et al., 2001; Koto et al., 2021; Kurfalı and Östling, 2021; Zeldes, 2017).
@@ A.4 Details of Probing Dataset Composition
Whenever possible, we rely on established probing datasets and transform instances into a unified format: 1) an input $x$ which is either one or a pair of span(s) or sentence(s), including the string and an optional starting and ending index in the context $c$ when task type is either a span or span-pair classification; 2) an optional textual context $c$ to encode $x$, for example the sentence in which a span occurs; and 3) a corresponding label $y$. If given, we use the original train/dev/test splits. However, if this division does not exist, we use a 70/10/20 ratio to form these splits. Furthermore, we adapt the design of some tasks to map to our task format. Exemplary, for the oLMmpics (Talmor et al., 2020) dataset, we transform the mask-filling tasks into a binary classification where the correct label corresponds to a sentence with a correctly filled mask and incorrect to a sentence where the mask was filled wrongly.
@@
Presented in Klafka and Ettinger (2020), we compose nine datasets to verify information about context words.
@@
Using the data presented in the BLiMP benchmark (Warstadt et al., 2020), we derive 67 probing datasets verifying specific phenomena, like island effect, covering morphology, syntax, and semantics. Unlike the original version, we compose a binary classification task for every phenomenon. Precisely, whether to accept or reject a given sentence, where rejecting means that the given linguistic phenomena is violated.
@@
For semantic relation classification we rely on the dataset of Hendrickx et al. (2010).
@@
The Decompositional Semantics Initiative [^6] https://decomp.io/ provides a large number of datasets to verify semantic phenomena. Apart of the common use semantic proto-roles (Rudinger et al., 2018a), we use their collection of works to compose probing datasets for factuality (Rudinger et al., 2018b), genericity (Govindarajan et al., 2019), event structure (Vashishtha et al., 2019), time (Vashishtha et al., 2019), and word sense (White et al., 2016).
@@
We form probing datasets addressing different lexical reasoning using the data presented in Talmor et al. (2020). As they provide multiple choices, we form correct instances by filling the gap with the correct option and wrong ones by filling in the other options. Specifically, we form dataset for always-never, age comparison, objects comparison, antonym-negation, multi-hop composition property conjunction, taxonomy conjunction, and encyclopedic composition.
@@
Phenomena
@@
anaphor agreement 3
determiner noun agreement 10
irregular forms 3
subject-verb agreement 10
Table 3: Overview of resources and linguistic phenomena mapping for morphology. It shows the number of datasets for the phenomena by dataset type.
Phenomena
@@
argument-structure 20
bigram-shift 1
binding 8
case 1
constituent parsing 2 1
control/raising 5
deoncausative-inchoative alternation 1
dependency parsing 1
ellipsis 3
filler-gap 9
island-effects 10
local attractor 1
object-number 2
part-of-speech 3
readability 1
sentence-length 1
subject-number 2
top-constituent-task 1
tree-depth 1
Table 4: Overview of resources and linguistic phenomena mapping for syntax. It shows the number of datasets for the phenomena by dataset type.
Phenomena
@@
Talmor et al. (2020)
age comparison 1
always-never 1
antonym negation 1
encyclopedic composition 1
multi-hop composition 1
negation 3 1 2 2
objects comparison 1
property conjunction 1
speculation 1 1 1
taxonomy connection 1
Table 5: Overview of resources and linguistic phenomena mapping for reasoning. It shows the number of datasets for the phenomena by dataset type.
Phenomena
@@
bridging 1 1
co-reference resolution 1
discourse connective 1
discourse representation theory 8
next-sentence prediction 1
rethorical structure theory 6 8
sentence order 1
Table 6: Overview of resources and linguistic phenomena mapping for discourse. It shows the number of datasets for the phenomena by dataset type.
Phenomena
@@
complex word identification 1
coordination inversion 1
event structure 4 2
factuality 1
genericity 6
metaphor 4
named-entity labeling 1
negative polarity item licensing 4
object-animacy 1
object-gender 1
passive 1
quantifiers 6
semantic relation classification 1
semantic proto-roles 20
semantic odd man out 1
semantic-role labeling 1
sentiment analysis 1
subject-animacy 1
subject-gender 1
synonym-/antonym-detection 1
tense 2
time 1
verb-dynamic 1
word content 1
word sense 1
Table 7: Overview of resources and linguistic phenomena mapping for semantics. It shows the number of datasets for the phenomena by dataset type.

### v4

@@ 6 Efficiency
Besides excluding licensed data (18 probing datasets), we analyze the effect of discarding training instances. As a result, we reduce the computation for encoding and the actual probing simultaneously. We follow Perlitz et al. (2023) and calculate the rank resolution, $95\%$ CI of model rank difference. This measure indicates the maximum expected rank deviation from evaluating an LM on FlashHolmes compared to Holmes. For example, a rank resolution of one means that an LM evaluated on FlashHolmes and Holmes has the same rank or switch place with its neighbors with a probability of 95%. Figure 13 shows the resulting rank resolution when training only on a fraction of the instances, from $1/2$ to $1/512$. Solely focusing on efficiency ($1/512$) still provides a decent rank resolution of ~2.6. In contrast, considering $1/2$ of the training data results in the best reliability of ~0.9. To balance benchmark reliability and efficiency, we compose FlashHolmes using $1/32$ of the training instances. Precisely, it reduces the computation expenses of evaluating LMs to ~3% of what Holmes would have required while preserving a high rank-correlation of ~1.5.
@@ 7 Related Work
Benchmarks approximate LMs abilities like general language understanding (Wang et al., 2019a, b), out-of-distribution generalization (Yang et al., 2023; Waldis et al., 2024b), real-world knowledge contradiction (Hou et al., 2024), adversarial scenarios (Nie et al., 2020; Wang et al., 2021), or retrieval (Thakur et al., 2021; Muennighoff et al., 2023). With the recent advent of large LMs, the predominant method has shifted to evaluate the obtained linguistic performance of LMs when providing textual instructions (Brown et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022). While LMs show substantial performance on application-oriented tasks (Liang et al., 2023) or mathematical reasoning (Cobbe et al., 2021), such evaluations are sensible to specific formulations (Mizrahi et al., 2024) or metrics (Schaeffer et al., 2023) employed. Thus, results of different benchmarks were found to disagree substantially (Yuan et al., 2024; Perlitz et al., 2024).
@@
Analyzing LMs’ linguistic competence started with static word vectors (Köhn, 2015), sentence embeddings (Conneau et al., 2018; Adi et al., 2017), the internals of translation models (Shi et al., 2016; Bau et al., 2019), or contextualized LMs (Tenney et al., 2019b, a; Hewitt and Manning, 2019). Other methodological work addressed the validity of obtained results with control tasks (Hewitt and Liang, 2019) or from an information theory perspective (Voita and Titov, 2020; Pimentel et al., 2020), or studied causal effects (Elazar et al., 2021). While further studies focus on whether LMs follow human understanding of linguistic competence when solving downstream tasks (Belinkov, 2022; Aw et al., 2023; Mahowald et al., 2024), Mosbach et al. (2020b) and Waldis et al. (2024a) found that downstream task fine-tuning hurts the understanding of linguistic phenomena.
In contrast to prior studies, Holmes assesses the linguistic competence of an extensive set of contemporary LMs covering a comprehensive collection of linguistic phenomena. Unlike other work evaluating linguistic phenomena (Blevins et al., 2023b; Amouyal et al., 2024) using prompting leading to unreliable results (Liang et al., 2023), probing allows Holmes to reliably and comprehensively compare LMs regardless of architecture or pre-training. As a result, Holmes can address recent calls to thoroughly and explicitly evaluate linguistic phenomena (Hu and Levy, 2023; Lu et al., 2023; Mahowald et al., 2024).
@@ 8 Conclusion
Holmes marks the most up-to-date and extensive consolidation of existing resources addressing the need to assess the linguistic competence of LMs in isolation. Our experiments demonstrate that LMs’ linguistic competence is pronounced regarding formal phenomena but lacks functional ones when information about broader textual contexts, such as rhetorical structure, is required. Simultaneously, size, architecture, and instruction tuning are crucial factors for differences among LMs. As LM and resources in linguistics constantly grow, we will actively extend Holmes with new datasets and upcoming LMs.
@@
Last Layer Internal Representation
Given the extensive scope of the analysis presented in this work, we focus on examining the internal representation of language models (LMs) through the output of their last layer. While this analysis provides valuable insights, it only partially captures the complexity inherent in LMs across all their layers. To facilitate further research into the comprehensive analysis of LMs, we see Holmes providing groundwork, including the release of the specific tasks in a unifying format and corresponding evaluation code, which can be easily adapted to investigate specific layers of LMs.
Coverage
We agree with Liang et al. (2023) and see one fundamental aspect in composing a benchmark in acknowledging its incompleteness. Linguistic phenomena, LMs, and underlying meta-studies are a subset of the variety of available resources. We consolidated them carefully to provide a comprehensive scope of the linguistic competence and various LMs. However, as benchmarks evolve as tools to assess LMs, we will further expand Holmes both with the existing and upcoming LMs and data resources.
@@
As Holmes relies on existing resources, it inherits the bias embodied in these datasets. Examples of such bias are gender equality or gender fairness, like the use of neo pronouns such as em in Lauscher et al. (2023).
Dataset Contamination
Holmes encompasses a large collection of established datasets, like OntoNotes (Weischedel et al., 2013). While we solely rely on LMs with open-sourced weights, the training or instruction-tuning data is not known for all of them, as for the Llama-2 (Touvron et al., 2023), Mixtral (Jiang et al., 2024), or Wizard (Xu et al., 2023) LMs. Therefore, we need to expect that some texts were part of the LMs’ pre-training corpora and that specific tasks, such as named-entity recognition (NER), were used during instruction tuning. However, instruction-tuning aligns LMs’ linguistic performance to produce coherent text responding to specific textual instruction provided and does not align LMs’ internal representations explicitly (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2024). As Holmes evaluates the linguistic competence using LMs’ internal representations, it retains its validity even under potential data contamination (Balloccu et al., 2024). Building upon our results, showing that downstream abilities are partly reflected in LMs’ internal representations, one could examine whether instruction-tuning injects task-specific information into LMs’ internal representations, thereby detecting task contamination.
@@ A.1 Additional Details on the Evolution of Probing Literature
We analyze publication trends by year and venue as shown in Table 3. Less work was published between 2015-2018 (earlier) focusing on LSTM-based (Linzen et al., 2016; Conneau et al., 2018) and static LMs (Köhn, 2015; Linzen et al., 2016; Belinkov et al., 2017; Conneau et al., 2018). With the release of BERT (Devlin et al., 2019) in 2019, we note increasing attention to analyzing linguistic abilities within LMs, with a peak of 90 papers in 2022. Considering the venue, more than half of the relevant work (149 papers) was published at major conferences (ACL and EMNLP), and 68 papers were published at AACL, EACL, NAACL, and COLING. [^5] Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-study. In addition, we observe a constant contribution of TACL, various workshops, such as Analyzing and Interpreting Neural Networks for NLP or Representation Learning for NLP.
@@ A.2 Experimental Details
A.3 Probing Datasets Categorization
We show in Table 4, Table 5, Table 8, Table 6, and Table 7 which resources Holmes use to cover morphology, syntax, semantics, reasoning, and discourse phenomena. Further, we provide illustrative examples of the phenomena. We rely on 33 works providing the data, the specific linguistic phenomena, or both. For example, for readability, we use the data of Weischedel et al. (2013) and calculated the Flesch score (Flesch, 1948).
@@ A.3 Probing Datasets Categorization
The second group of 75 tasks verifies the following syntax phenomena: Part-of-speech and constituent labeling (Weischedel et al., 2013); dependency labeling (Silveira et al., 2014); bigram-shift (whether two words were shifted), tree-depth (the depth of a sentence constituency tree), top-constituent-task (top constituency tag), and sentence-length (Conneau et al., 2018); subject- & object-number (singular/plural), and deoncausative-inchoative alternation (interaction of a verb with its context) based on Klafka and Ettinger (2020); binding, control/raising, negative polarity item licensing, island-effects, argument-structure, ellipsis, and filler-gap (Warstadt et al., 2020; Huebner et al., 2021).
@@
Third, consider 67 datasets covering semantics phenomena: Named-entity labeling and semantic-role labeling (Weischedel et al., 2013); tense, semantic odd man out, word content, and coordination inversion (Conneau et al., 2018); semantic relation classification (Hendrickx et al., 2010); semantic proto-roles (Rudinger et al., 2018a); factuality (if a span is factual or not) (Rudinger et al., 2018b); genericity (whether a span is generic or not) (Govindarajan et al., 2019); event structure (Gantt et al., 2022); time (time dimension of a span) (Vashishtha et al., 2019); word sense (White et al., 2016); sentiment analysis (Socher et al., 2013); object- and subject-animacy (whether a entity is animate, like human, or not, such as cars), object- and subject-gender (male/female), verb-tense, and verb-dynamic Klafka and Ettinger (2020); metaphor (Mohler et al., 2016; Birke and Sarkar, 2006; Steen et al., 2010); complex word identification (whether the word is complex or not) (Paetzold and Specia, 2016); and passive (Krasnowska-Kieraś and Wróblewska, 2019). In addition, we derive an synonym-/antonym-detection dataset using WordNet (Miller, 1995) and the texts from OntoNotes v5 Weischedel et al. (2013).
@@
Forth, 19 datasets cover reasoning phenomena: Paraphrasticity with negation and antonyms (Vahtola et al., 2022); negation detection (Szarvas et al., 2008; Konstantinova et al., 2012; Morante and Blanco, 2012); negation-span classification (does a span cause a negation) (Szarvas et al., 2008; Konstantinova et al., 2012); negation-correspondence (the target span of a negation) (Szarvas et al., 2008; Konstantinova et al., 2012); speculation detection, speculation-span classification, and speculation-correspondence (the target span of a sepculation) (Szarvas et al., 2008); and always-never, age comparison, objects comparison, antonym negation, property conjunction, taxonomy connection, and multi-hop composition (Talmor et al., 2020b).
@@
Finally, Holmes embodies 28 datasets addressing discourse phenomena: Co-reference resolution Weischedel et al. (2013); bridging (Hou, 2018, 2020; Pandit and Hou, 2021); discourse connective (Nie et al., 2019); sentence order and next-sentence prediction (Narayan et al., 2018); Given discourse tree, whether two nodes correspond (discourse correspondence), the correct order of two nodes (discourse order), node-node relation (discourse relation), distance between two nodes (discourse distance), explicit node class discourse explicit classes, implicit node class discourse implicit classes (Webber et al., 2019; Kurfalı and Östling, 2021); and given a rhetorical tree with the number of child nodes (rst-count), the node depth (rst-depth), distance between two nodes rst-distance, node-node relation (rst-relation), node-node relation group (rst-relation-group), appear two nodes after each other (rst-successively), node type (rst-type) (Carlson et al., 2001; Koto et al., 2021; Kurfalı and Östling, 2021; Zeldes, 2017).
@@ A.4 Details of Probing Dataset Composition
Whenever possible, we rely on established probing datasets and transform instances into a unified format: 1) an input $x$ which is either one or a pair of span(s) or sentence(s), including the string and an optional starting and ending index in the context $c$ when task type is either a span or span-pair classification; 2) an optional textual context $c$ to encode $x$, for example the sentence in which a span occurs; and 3) a corresponding label $y$. Figure 14 shows the composition of the specific probing input $x$ for these four tasks using the internal representation of the last layer of LMs. Note that additional averaging operations are required if words are tokenized into multiple tokens to get one average vector representing one word, for example, when probing for the part-of-speech tag of a rare word.
Figure 14: Overview of the composition of the probing input $x$ based on the given text the four types of tasks using concatenating and averaging. In case a tokenizer splits one word into multiple tokens and applies additional averaging operations, such as when probing the part-of-speech phenomenon.
If given, we use the original train/dev/test splits. However, if this division does not exist, we use a 70/10/20 ratio to form these splits. Furthermore, we adapted the design of some data to map our dataset format. Exemplary, for the oLMmpics (Talmor et al., 2020b) dataset, we transform the mask-filling tasks into a binary classification where the correct label corresponds to a sentence with a correctly filled mask and incorrect to a sentence where the mask was filled wrongly.
@@
Presented in Klafka and Ettinger (2020), we compose the nine datasets verifying LMs’ knowledge about the context of words. For example, is a word animate (like animals or humans) or inanimate (like buildings or vehicles), or is a verb static or dynamic.
@@
Using the data presented in the BLiMP benchmark (Warstadt et al., 2020), we derive 67 probing datasets verifying specific phenomena, like island effect, covering morphology, syntax, and semantics. Unlike the original version, we compose a binary classification task for every phenomenon, either accepting a valid sentence or rejecting one that violates the given linguistic phenomenon.
@@
For semantic relation classification, we rely on the dataset of Hendrickx et al. (2010).
@@
The Decompositional Semantics Initiative [^6] https://decomp.io/ provides a large number of datasets to verify semantic phenomena. Apart from the common use semantic proto-roles (Rudinger et al., 2018a), we use their collection of works to compose probing datasets for factuality (Rudinger et al., 2018b), genericity (Govindarajan et al., 2019), event structure (Vashishtha et al., 2019), time (Vashishtha et al., 2019), and word sense (White et al., 2016).
@@
We form probing datasets addressing different lexical reasoning using the data presented in Talmor et al. (2020b). As they provide multiple choices, we form correct instances by filling the gap with the correct option and wrong ones by filling in the other options. Specifically, we form dataset for always-never, age comparison, objects comparison, antonym-negation, multi-hop composition property conjunction, taxonomy conjunction, and encyclopedic composition.
@@
Figure 15: Detailed Holmes vs. HELM (Liang et al., 2023) comparison for 40 open decoder models and 22 Blimp datasets covering quantifier, island effects, irregular forms, and binding phenomena. We use the evaluation code of HELM and run the prompting-based adaption (multiple joice joined). The Holmes and Helm results for 40 open decoder models. These results show the advantage of disentangled evaluation (Holmes) over entanglement evaluations (like in HELM), which intertwine the understanding of specific linguistic phenomena and other abilities (like following instructions or answering precisely) in HELM. Most HELM results are below the random baseline, underscoring the necessity to measure linguistic phenomena directly in isolation within LMs.
Phenomena Illustrative Example
@@
anaphor agreement Katherine can’t help herself*/himself. 3
determiner noun agreement Craig explored that grocery store*/stores. 10
irregular forms Edward hid*/hidden the cats. 3
subject-verb agreement A sketch of lights does not*/do not appear. 10
Table 4: Overview of resources and linguistic phenomena mapping for morphology. We give an illustrative example for each phenomenon (*indicates the right option, if options are given) and the number of datasets for the phenomenon by dataset type.
Phenomena Illustrative Example
@@
argument-structure Most cashiers are disliked*/flirted. 20
bigram-shift What are you*/you are doing out there? 1
binding Carlos said that Lori helped him*/himself. 8
case-subjective-pronoun He brought the pig this suit.*/The pig brought he this suit. 1
constituent parsing sees Bill $\Rightarrow$ VP 2 1
control/raising Julia wasn’t fun*/unlikely to talk to. 5
deoncausative-inchoative alternation The warden melted the ice.*/The warden bought the ice. 1
dependency parsing (into, air) $\Rightarrow$ pobj 1
ellipsis He cleans one important book and Stacey cleans a few.*/He cleans one book and Stacey cleans a few important. 3
filler-gap Brett knew what many waiters find.*/Brett knew that many waiters find. 9
island-effects Which bikes is John fixing?*/Which is John fixing bikes? 10
local attractor Can the access work?*/ Can the access works? 1
object-number Oh gods! $\Rightarrow$ Plural 2
part-of-speech cucumber $\Rightarrow$ NN (Noun Singular) 3
readability Curriculums need selling points. $\Rightarrow$ 50.5 (middle) 1
sentence-length Oh gods! $\Rightarrow$ 3 words 1
subject-number Things are going to be noticed. $Rightarrow$ Plural 2
top-constituent Did it all matter? $Rightarrow$ VBD NP VP 1
tree-depth Where do you want it? $\Rightarrow$ 6 1
Table 5: Overview of resources and linguistic phenomena mapping for syntax. We give an illustrative example for each phenomenon (*indicates the right option, if options are given) and the number of datasets for the phenomenon by dataset type.
Phenomena Illustrative Example
@@
Talmor et al. (2020b)
age comparison 21 years old is older than 35 years fold.*/21 years old is younger than 35 years fold. 1
always-never Horses have always*/never four legs. 1
antonym negation It was not*/really hot, it was cold. 1
multi-hop composition Comparing a 23, a 38 and a 31 year old, the last*/first is oldest. 1
negation I don’t like bananas. $\Rightarrow$ Negation 3 1 2 2
objects comparison An airplane is bigger*/smaller than a pen. 1
property conjunction A pen*/computer is usually located at hand and used for writing. 1
speculation Just about every PC can be upgraded. $\Rightarrow$ Speculation 1 1 1
taxonomy connection Ferry and floatplane are both boats*/airplaines. 1
Table 6: Overview of resources and linguistic phenomena mapping for reasoning. We give an illustrative example for each phenomenon (*indicates the right option, if options are given) and the number of datasets for the phenomenon by dataset type.
Phenomena Illustrative Example
@@
bridging The disease and symptoms of advanced infection. $\Rightarrow$ Valid Bridge 1 1
co-reference resolution National Taiwan University opened the doors of five of its graduate schools. $\Rightarrow$ Valid Co-Reference 1
discourse connective Leaning against his hip. He reclined with his feet up on the table. $\Rightarrow$ when 1
discourse representation theory This is an old story. We’re talking about years ago. $\Rightarrow$ Implicit Relation 8
next-sentence prediction Sentence A, Sentence B $\Rightarrow$ Valid Next Sentence 1
rethorical structure theory The statistics quoted by the " new " Census Bureau report $\Rightarrow$ Elaboration 6 8
sentence order Given Sentence B, C, and D $\Rightarrow$ C is at position 2 1
Table 7: Overview of resources and linguistic phenomena mapping for discourse. We give an illustrative example for each phenomenon (*indicates the right option, if options are given) and the number of datasets for the phenomenon by dataset type.
Phenomena Illustrative Example
@@
complex word identification membrane $\Rightarrow$ Complex, his $\Rightarrow$ Simple 1
coordination inversion He knew it, and he deserved no answer. $\Rightarrow$ Inversion 1
event structure Give them to a library or burn them. $\Rightarrow$ Distributive 4 2
factuality I ran across this item on the Internet. $\Rightarrow$ Factual 1
genericity I assume you mean the crazy horse memorial. $\Rightarrow$ Not Dynamic 6
metaphor After all, morons pay taxes, too. $\Rightarrow$ Valid Metaphor 4
named-entity labeling Paris $\Rightarrow$ City 1
negative polarity item licensing Only/Even Bill would ever complain. 4
object-animacy The rhino fined the pumpkin. $\Rightarrow$ Animate 1
object-gender The princess uncovered the heiress. $\Rightarrow$ Feminine 1
passive He is considered a European poet through and through. $\Rightarrow$ Passive Sentence 1
quantifiers There aren’t many*/all lights darkening. 6
semantic relation classification Those cancers were caused by radiation exposures. $\Rightarrow$ Cause-Effect 1
semantic proto-roles These look fine to me. $\Rightarrow$ Exists as physical 20
semantic odd man out I wanted to know if it was real or a ploy. $\Rightarrow$ Original 1
semantic-role labeling And what effect does their return have on campus? $\Rightarrow$ ARGM-ADV 1
sentiment analysis You ’ll probably love it. $\Rightarrow$ Positive 1
subject-animacy The turtle betrayed the judge. $\Rightarrow$ Animate 1
subject-gender The waitress betrayed the judge. $\Rightarrow$ Feminine 1
synonym-/antonym-detection Is the degree really that important $\rightarrow$ unimportant to them? $\Rightarrow$ Antonym Replacement 1
tense I quietly snuck up to him and pulled at his sleeve. $\rightarrow$ Present 2
time His mother was also killed in the attack. $\Rightarrow$ Minutes 1
verb-dynamic The lawyer found the judge. $\Rightarrow$ Dynamic Verb 1
word content You mean Alice. $\Rightarrow$ Contains Word Alice 1
word sense His mother was also killed in the attack. $\Rightarrow$ Supersense Noun Person 1
Table 8: Overview of resources and linguistic phenomena mapping for semantics. We give an illustrative example for each phenomenon (*indicates the right option, if options are given) and the number of datasets for the phenomenon by dataset type.